The normal to the parabola y2= 4ax at the point P(at2 , 2at) meets the x-axis at A. Find the equation of the locus of the midpoint of AP as t varies.


1
Expert's answer
2021-08-25T18:33:22-0400

To find the normal;

At A(y,x) but y=0

Since;


y2=4ax

"2y \frac{dy}{dx} =4a"

"\frac{dy}{dx} = \frac{2a}{y}"

Slope of the tangent at P(at2 , 2at); "\frac{dy}{dx} =\frac{2a}{2at} =\frac{a}{t}"


Hence the slope at normal P is given as "\frac{-1}{\frac{1}{t}} =-t"

Therefore equation of the normal at P is given by "tx+y=2at+at^3"

Hence;

"y-2at=-t(x-at^2)"

But y=0

"−2at=−tx+at^3 \\ 
 

Rewrite \,as \\

tx=at^3+2at\\

Divide \,by \,t,we \,get\\

x=at^2+2ax=at^2
 +2a"


Take the midpoint of AP to be B(h,k)

Here


"h=\frac{at^2 +2a +at^2}{2} =at^2 +a"

"k= \frac{0+2at}{2}= at"


since; y "^2 =4ax"

We have the equation of the locus as t varies given by

"(2at)^2 =4a(at^2)"

Using h and k ,we can make the following substitutions


"(2 x k)^2 =4a(h-a )"

"4k^2 =4ah-4a^2"

Rewrite as

"k^2 =ah -a^2"

"k^2 - ah +a^2 =0"

Write in terms of x and y as


"y^2-ax -a^2=0"